OpenStudy (anonymous):

(Will give MEDAL) Which is the equation of the given line in point-slope form? Coordinates: (1/2, 3) and (-4, -1)

OpenStudy (anonymous):

y=8/9+2.5555555555

OpenStudy (anonymous):

Answer choices are: A. y + 4 = 9/8 (x + 1) B. y + 1 = 8/9 (x + 4) C. y + 1 = 8/7 (x + 4) D. y - 1 = 8/9 (x - 4) So that can't be it.